OpenStudy (anonymous):

Which function's graph has a period of 2? A. y=cos( x-pi/2) B. y= -4sin2x c. y= 2 sinpix D. y=3cosx

OpenStudy (anonymous):

period of \(\sin(bx)\) is \(\frac{2\pi}{b}\)

OpenStudy (anonymous):

set \(\frac{2\pi}{b}=2\) solve for \(b\)

OpenStudy (anonymous):

you should get \(b=\pi\) so you are looking for \(\sin(\pi x)\)

OpenStudy (anonymous):

so is C?

OpenStudy (anonymous):

yes